OpenStudy (anonymous):

x/9 = 4 4/9 9/4 1/36 36

undeadknight26 (undeadknight26):

agreed.

OpenStudy (anonymous):

Solve. 2(p + 1) = 18 (Points : 4) 8.5 9.5 8 10

OpenStudy (jojo4eva):

8

OpenStudy (jojo4eva):

PEMDAS. 8+1=9, 2*9=18

OpenStudy (jojo4eva):

i love math so i understand a lot about it lol

OpenStudy (anonymous):

-1/5x = 6 -30 -6/5 -1 6 1/5

OpenStudy (jojo4eva):

juss kee[ feedin them to me and ill help u with them. just mention me in it

OpenStudy (jojo4eva):

-30

OpenStudy (anonymous):

yay so i was right

OpenStudy (jojo4eva):

its a positive 6 so x has to be negative

OpenStudy (anonymous):

By what number must you multiply each side of the equation 4/5x =8 to produce the equivalent equation x = 10 5/4 5 1/5 -4/5

OpenStudy (anonymous):

at this point im your biggest fan i just really wish i loved math like you so i could understand it even more and not be worried about anything so i wanna say thank you @jojo4eva

OpenStudy (jojo4eva):

no problems anytime u need help let me kno and let me see about this question u juss posted

OpenStudy (anonymous):

ok

OpenStudy (jojo4eva):

its 5/4

OpenStudy (jojo4eva):

ur multiplying by the reciprocal of 4/5

OpenStudy (jojo4eva):

therefore its 5/4

OpenStudy (jojo4eva):

anything else?

OpenStudy (anonymous):

Solve. 8x = 4x – 32 (Points : 1) 8 4 –4 –8

OpenStudy (jojo4eva):

Hey i gotta help you through this one i can just give you the answer so im gonna help you through it ok?

OpenStudy (jojo4eva):

Now what you wanna do is compare like terms so you have to put the same variable on the same side so tell me what you should do

OpenStudy (anonymous):

-8

OpenStudy (jojo4eva):

BOOYYAA you got it

OpenStudy (jojo4eva):

Good jobb

OpenStudy (anonymous):

thank u

OpenStudy (jojo4eva):

no probs if you need any more help just mention me in it and ill be there